# Evolutionary Optimization of Physics-Informed Neural Networks: Survey and Prospects

### v2

@@ V Conclusion
However, like any new technique, PINNs inevitably call for new technical innovations. Many of the approaches commonly used in training standard data-driven neural nets need to be re-evaluated, such as the method of network initialization and the choice of activation functions. The use of gradient-based optimization alone, which includes the hugely successful family of SGD algorithms, is especially worthy of reconsideration in the light of analyses summarized in this review. Identified challenges pertaining to local minima traps and vanishing gradients of physics-informed learning objectives highlight how these can prove deleterious for PINNs. These obstacles to conventional gradient-based optimization serve to outline ways in which gradient-free algorithms of computational intelligence may be uniquely advantageous as substitutes to or in synergy with SGD. Indeed, the crafting of synergistic optimization frameworks that meld gradient-based and gradient-free optimization unveils novel research questions and directions, firmly positioning scientific machine learning as an emerging area of translational interest for the computational intelligence community.
In fact, EAs have demonstrated their potential not only in PINNs but also in various approaches to solving PDEs, such as genetic programming [170, 171, 172] and reduced order models [173]. Genetic programming leverages symbolic regression to derive solutions to PDEs, offering the distinct advantage of high interpretability in the resulting solutions [174]. This capability underscores its potential for advancing the development of interpretable PDE solvers, thereby fostering enhanced understanding and usability in scientific and engineering applications. Similarly, evolution strategies have proven effective in improving the prediction accuracy of reduced order models, driven by their robust global search capabilities. The aforementioned cases highlight the significant potential of evolutionary computation in solving PDEs. We believe that this survey can serve as a valuable resource for guiding future advancements in this promising research direction.

### v3

@@ IV-D3 Meta-Learning Neural Architecture, Loss Function, and Optimizer
IV-D4 Other Approaches for Advancing Generalizability
A notable alternative to generalized PINNs is through neural operator networks [194, 195, 196], which, by learning mappings between infinite-dimensional function spaces, allow models to solve PDEs in real-time under many different ICs or BCs, requiring only light task-specific fine-tuning. However, unlike parameterization-agnostic Baldwinian-PINNs, neural operators call for explicit parameterization of the conditions over which the model is to generalize. Although this does make neural operators less flexible compared to meta-learning, they are considered an important complementary approach to model generalization. At the time of writing of this paper, the effectiveness of EAs for optimizing physics-informed neural operator nets remains an open question. Some foundational work bringing randomization into shallow operator networks has appeared in [197], showing their potential to outperform standard operator nets by several orders of magnitude in terms of approximation accuracy and computational cost.
IV-E Evolving Interpretable Models Beyond PINNs
In scientific machine learning, a key objective of discovering new insights from experimental data relies on human-interpretable models or outputs. While advances have been made in explaining the outputs of deep neural networks, these models are largely perceived as black-boxes [198]. The same issue naturally afflicts PINNs, prompting the brief discussion below on the evolution of other model types.
In contrast to neural nets, the subfield of genetic programming in evolutionary computation has a long history of successfully discovering free-form natural laws from data through symbolic regression [199, 200]. Recently, these algorithms have even been used for solving high-dimensional PDEs, with the distinct advantage of offering compact analytical expressions approximating complex solutions [201, 202, 203]. Balancing compactness and model accuracy is indeed a non-trivial task, one that is well-suited to evolutionary multi-objective optimization. In addition to symbolic regression, a highly interpretable mathematical description of dynamical systems can be obtained by identifying a best-fit linear dynamics model (i.e., a linear system of governing ODEs) from data. Although finding such a data-driven model that simultaneously satisfies physics constraints translates into a highly constrained, nonlinear and non-convex optimization problem [204], it has been shown to be efficiently solvable by evolution strategies [205]. Adaptations of this evolutionary approach to identify nonlinear dynamics, through neural ODEs [206], have also been explored [207], with the final prediction accuracy and training stability of evolution exceeding those of models trained by gradient descent.
@@ V Conclusion
However, like any new technique, PINNs inevitably call for new technical innovations. Many of the approaches commonly used in training standard data-driven neural nets need to be re-evaluated, such as the method of network initialization and the choice of activation functions. The use of purely gradient-based optimization, which includes the hugely successful family of SGD algorithms, is especially worthy of reconsideration in the light of analyses summarized in this work. Identified challenges pertaining to local minima traps, fixed point traps, and vanishing gradients of physics-informed learning objectives highlight how these can prove deleterious for PINNs. These obstacles to conventional gradient-based optimization serve to outline ways in which gradient-free algorithms of computational intelligence may be uniquely advantageous as substitutes to or in synergy with SGD. Indeed, the crafting of synergistic optimization frameworks that meld gradient-based and gradient-free optimization unveils novel research questions and directions, firmly positioning scientific machine learning as an emerging area of translational interest for the computational intelligence community.